OpenStudy (anonymous):

the product of 9 and x is the same as2 subtracted from the product of 7and x ,find x

OpenStudy (anonymous):

9x = 7x-2 Do you understand how to get that?

OpenStudy (anonymous):

so what is x do I sub or add 9x,7x?

OpenStudy (anonymous):

help with what is x

OpenStudy (anonymous):

subtract 7x from both sides and see what happens

OpenStudy (anonymous):

You want your x on one side. You can either sub 9 or sub 7

OpenStudy (anonymous):

ok so I have 2 as answer right 9x-7x=2and7x-7x=0 so is x 2read the problem to see if correct please

OpenStudy (anonymous):

the product of 9 and 2 is the same as2 subtracted from the product of 7and 2 ?

OpenStudy (anonymous):

18 = 12

OpenStudy (anonymous):

what I'm more confused please how did you get this? this is how much i did9x-7x 7x7x 2x

OpenStudy (anonymous):

the product of 9 and X is the same as2 subtracted from the product of 7and X 9x = 7x - 2 subtract 7x -7x -7x --------------- 2x = 0x -2

OpenStudy (anonymous):

I'll make it really simple. 9x = 7x-2 You need your x on one side. So think of bringing the 7x from the right side to the left side, but remember when you bring it over, it becomes -7x. Whenever you bring something over from one side to another, change the sign. Therefore: 9x - 7x = -2 Which is: 2x = -2 Now divide both sides by 2 x = -1

OpenStudy (anonymous):

Maybe that's easier to understand than subtracting this or that from both sides or whatnot.

OpenStudy (anonymous):

Understand?

OpenStudy (anonymous):

Petewe!!!!!!! wow you made it seem simple I am so glad to have found this website THANK-YOU<YOU DON'T know how helpful you have been I have one last problem to post

OpenStudy (anonymous):

No problem, message me if you need anything.